When 24.7 cal of heat is added to 18.8 g of aluminum at 22.0 ∘C, what is the final temperature of the aluminum?
1
Expert's answer
2017-09-06T10:36:06-0400
Q=C∙m∙∆T
Specific heat capacity of Aluminum:C= 920 J/(kg∙℃)
1 cal=4.187 J,
24.7-x J; x=24.7∙4.187=103.42 J
103.42 J=920 J/(kg∙℃)∙0.0188 kg∙∆T
∆T=103.42/(920∙0.0188)=6℃
∆T=T_2-T_1,where T_2=∆T+T_1=6+22=28℃.
Answer: 28℃.